\begin{thebibliography}{99}

\footnotesize{
\bibitem{snyder}  H. Snyder, Phys. Rev. {\bf 71}, 38 (1947); ibid. {\bf 72},
68 (1947).

\bibitem{noncom}  S. Doplicher, K. Fredenhagen and J. E.Roberts, Phys. Lett. 
B{\bf 331}, 39 (1994); S. Majid and H. Ruegg, Phys. Lett. B{\bf 334}, 348 (1994); J.
Lukierski, H. Ruegg and W. J. Zakrzewski, Ann. Phys. {\bf 243}, 90 (1995);
N. Seiberg and E. Witten, JHEP {\bf 9909}, 032 (1999), hep-th/9908142; I. F.
Riad and M. M. Sheikh-Jabbari, JHEP {\bf 08}, 045 (2000), hep-th/0008132; M.
M. Sheikh-Jabbari, Phys. Rev. Lett. {\bf 84}, 5265 (2000), hep-th/0001167;
J. Madore, S. Schraml, P. Schupp and J. Wess, Eur. Phys. J. C{\bf 16}, 161 (2000).

\bibitem{kruglov}  S. I. Kruglov, Ann. Fond. Broglie {\bf 27}, 343 (2002), hep-th/0110059; B.
G. Sidharth, Ann. Fond. Broglie {\bf 27}, 333 (2002), physics/0110040.

\bibitem{FD}  F. Dyson, Am. J. Phys. {\bf 58}, 209 (1990). S. Tanimura, Ann.
Phys. {\bf 220}, 229 (1992); M. Land, N. Shnerb and L. Horwitz,
hep-th/9308003.

\bibitem{brown}  K. R. Brownstein, Am. J. Phys. {\bf 67}, 639 (1999).

\bibitem{landau}  L. D. Landau and E. M. Lifshitz, {\it The Classical Theory
of Fields.} 4th revised English ed. [Translated from the 6th Russian
edition] (Pergamon Press, 1979). 402p., \S 63.

\bibitem{ivezic} L. \v{S}kovrlj and T. Ivezi\'{c}, hep-ph/0203116.

\bibitem{jackson}  J. D. Jackson, hep-ph/0203076.

\bibitem{hb}  M. Jacob and G. C. Wick, Ann. Phys. {\bf 7}, 404 (1959); H. M.
Ruck and W. Greiner, J. Phys. G: Nucl. Phys. {\bf 3}, 657 (1977).

\bibitem{elmnoncom} A. Connes, M. R. Douglas, and A. Schwarz, JHEP 02, 003 (1998), hep-th/9711162; 
D. Galtsov and V. Dyadichev, hep-th/0012059.

\bibitem{prl} S. M. Carroll, J. A. Harvey, V. A. Kostelecký, C. D. Lane and 
T. Okamoto, Phys. Rev. Lett. {\bf 87}, 141601 (2001).

\bibitem{kadysh}  V. G. Kadyshevsky, Nucl. Phys. B{\bf 141}, 477 (1978); V.
G. Kadyshevsky, M. D. Mateev, R. M. Mir-Kasimov and I. P. Volobuev, Theor.
Math. Phys. {\bf 40}, 800 (1979) [Teor. Mat. Fiz. {\bf 40}, 363 (1979)]; V.
G. Kadyshevsky and M. D. Mateev, Phys. Lett. B{\bf 106}, 139 (1981); V. G.
Kadyshevsky and M. D. Mateev, Nuovo Cim. A{\bf 87}, 324 (1985); A. D.
Donkov, R. M. Ibadov, V. G. Kadyshevsky, M. D. Mateev and M. V. Chizhov,
ibid. {\bf 87}, 350 (1985); ibid. {\bf 87}, 373 (1985). See also the old
work: V. G. Kadyshevsky, R. M. Mir-Kasimov and N. B. Skachkov, Yad. Fiz. 
{\bf 9}, 212 (1969). See also references to the old works of M. Markov in
the above-cited papers.

\bibitem{amelino}  A. Kempf, G. Mangano and R. B. Mann, Phys. Rev. D{\bf 52}%
, 1108 (1995); G. Amelino-Camelia, Nature {\bf 408}, 661 (2000);
gr-qc/0012051; hep-th/0012238; gr-qc/0106004; J. Kowalski-Glikman,
hep-th/0102098; G. Amelino-Camelia and M. Arzano, hep-th/0105120; N. R.
Bruno, G. Amelino-Camelia and J. Kowalski-Glikman, hep-th/0107039.

\bibitem{dvoeglazovaps} V. V. Dvoeglazov, AIP Conf. Proc. {\bf 1191}, 80 (2009).
}

\newpage

\begin{center}

\title{{\bf NOTE ADDED:\\
Question on Iterated  Limits in Relativity}\\}         % Enter your title between curly braces
\author{V. V. Dvoeglazov\\UAF, Universidad de Zacatecas\\
Apartado Postal 636, Suc. 3\\ Zacatecas 98061 Zacatecas, M\'exico}        % Enter your name between curly braces
\date{\empty}          % Enter your date or \today between curly braces
%\maketitle

\end{center}

\begin{abstract}
ABSTRACT. Two iterated limits are not equal each other, in general. Thus, we present an example when the massless limit of the function of $E, {\bf p}, m$ does not exist in some calculations within quantum field theory.

RESUM\'E. Deux limites it\'er\'ees ne sont pas \'egales en g\'en\'eral. Ainsi, nous pr\'esentons un exemple o\'u la limite sans masse de la fonction de $E, {\bf p}, m$ n'existe pas dans certains calculs de la th\'eorie quantique des champs.

KEYWORDS: Iterated limits, Relativity, Massless limit.
\end{abstract}

\date{\empty}
%\maketitle

%\vspace{10mm}

In the previous paper~\cite{VVDPE} we found some intrinsic contradictions related to the mathematical foundations of modern physics.
It is well known that the partial derivatives commute in the Minkowski space (as well as in the 4-dimensional momentum space).
However, if we consider that the energy is the implicit function of the 3-momenta and mass (thus, approaching the mass hyperboloid formalism,
$E^2 - {\bf p}^2 c^2 = m^2 c^4$)
then we may be interested in the commutators of  the whole-partial derivatives~\cite{Brown} instead. The whole-partial derivatives do not commute, in general. If they are associated with the corresponding physical operators, we would have the uncertainty relations for dynamically-conjugated physical quantities in the latter case.
This is an intrinsic contradiction. While we start from the same postulates, on using two different ways of reasoning we arrive at the two different physical conclusions.

In the present note I would like to ask another question related to the mathematical foundations of the special relativity.  Sometimes, when calculating dynamical invariants (and other physical quantities in quantum field theory), and when studying the corresponding massless limits we need to calculate the iterated limits. We may encounter a rare situation when two iterated limits are not equal each other in physics!? See, for example, Ref.~\cite{VVD1}.
We were puzzled calculating the iterated limits of the aggregate $\frac{E^2 - {\bf p}^2}{m^2}$ (or the inverse one, $\frac{m^2}{E^2 -{\bf p}^2}$, \,\, $c=1$).\footnote{Similar mathematical examples are  presented in  https://en.wikipedia.org/wiki/Iterated\,\,\,limit.}
%\vspace{-5mm}
\begin{eqnarray}
\lim\limits_{m\to 0} \lim\limits_{E\to \pm\sqrt{{\bf p}^2 +m^2}} (\frac{m^2}{E^2 -{\bf p}^2}) &=& 1\,,\\
\lim\limits_{E\to \pm\sqrt{{\bf p}^2 +m^2}} \lim\limits_{m\to 0} (\frac{m^2}{E^2 -{\bf p}^2}) &=& 0\,.
\end{eqnarray}
Physics should have the well-defined dynamical invariants. Which iterated limit should be applied in the study of massless limits?

The question of the iterated limits is studied in~\cite{Ilyin1,Ilyin2}. However, the answers
leave the room for misunderstandings and contradictions with the experiments.

I am grateful to the referee for his clear reports. I am grateful to R. Keys for his help with the French translation.

%\vspace{-5mm}

\begin{thebibliography}{99}

%\footnotesize{

\bibitem{VVDPE} V. V. Dvoeglazov, Phys. Essays, this issue.\\[-6mm]

\bibitem{Brown} K. R. Brownstein, Am. J. Phys. {\bf 67}, 639 (1999).\\[-6mm]

\bibitem{VVD1} V. V. Dvoeglazov. Int. J. Theor. Phys. {\bf 37}, 1915 (1998);  Czech.  J.  Phys. {\bf 50}, 225 (2000).\\[-6mm]

\bibitem{Ilyin1} V. A. Ilyin, V. A. Sadovnichii and  B. Kh. Sendov, {\it Mathematical Analysis. Continuation.} (MSU, Moscow, 1987), in Russian.\\[-6mm]

\bibitem{Ilyin2}  V. A. Ilyin and E. G. Poznyak, {\it Fundamentals of Mathematical Analysis.} Vol. 1. Ch. 14. Functions of Several Variables
(Fizmatlit, Moscow, 2002), in Russian.\\[-6mm]

%}

\end{thebibliography}

\documentstyle[12pt]{article}

\begin{document}
\title{A Note on Non-commutativity}
\author{{\bf Valeri V. Dvoeglazov}\\
Universidad de Zacatecas\\
Apartado Postal 636, Suc. 3\\
Zacatecas 98061 Zacatecas, M\'exico\\
E-mail: valeri@fisica.uaz.edu.mx}

\date{\empty}
\maketitle


\begin{abstract}
ABSTRACT. Ambiguities have recently been found in the definition of the partial
derivative (in the case of presence of both explicit and implicit
dependencies of the function subjected to differentiation). 
We investigate the possible influence of this subject on quantum
mechanics and the classical/quantum field theory. Surprisingly, some
commutators of operators of space-time 4-coordinates and those of 4-momenta 
are {\it not} equal to zero.

RESUM\'E. Des ambigu\"it\'es ont \'et\'e r\'ecemment trouv\'ees dans la d\'efinition de la d\'eriv\'ee partielle (dans le cas de pr\'esence de d\'ependances \`a la fois explicites et implicites de la fonction soumise \`a la diff\'erenciation). Nous \'etudions l'influence possible de ce sujet sur la m\'ecanique quantique et la th\'eorie classique / quantique des champs. Fait surprenant, certains commutateurs 
d'op\'erateurs de coordonn\'ees spatio-temporelles et ceux de 4-moments ne sont pas \'egaux \`a z\'ero.

KEYWORDS: Non-commutativity, quantum mechanics, whole-partial derivatives 

PACS: 04.62.+v    02.40.Gh    02.30.-f
\end{abstract}

The assumption that the operators of
coordinates do {\it not} commute $[\hat{x}_{\mu },\hat{x}_{\nu }]\neq 0$
has been made by H. Snyder~\cite{snyder}. The Lorentz symmetry thus may be
broken. This idea~\cite
{noncom,kruglov} received attention in the context of ``brane theories''.
Moreover, the famous Feynman-Dyson proof of Maxwell equations~\cite{FD}
contains intrinsically the non-commutativity of velocities $[\dot x_i (t), 
\dot x_j (t)] \neq 0$ that also may be considered as a contradiction with
the well-accepted theories (while there is no any contradiction therein).

On the other hand, it was recently discovered that the concept of partial
derivative is {\it not} well defined in the case of both explicit and
implicit dependence of the corresponding function, which the derivatives act
upon~\cite{brown}.
The
well-known example of such a situation is the field of an accelerated
charge~ \cite{landau}.\footnote{%
Firstly, Landau and Lifshitz wrote that the functions depended on $t^{\prime }$,
and only through $t^{\prime }+R(t^{\prime })/c=t$ they depended implicitly
on $x,y,z,t$. However, later (in calculating the formula (63.7)) they used
the explicit dependence of $R$ on the space coordinates of the
observation point too. 
Jackson~\cite{jackson} agrees with~\cite{landau} that one should find ``a contribution to the
spatial partial derivative for fixed time $t$ from explicit spatial
coordinate dependence (of the observation point).''} 
\v{S}kovrlj and Ivezi\'{c}~\cite{ivezic} call this partial derivative as `{\it complete} partial
derivative'; Chubykalo and Vlayev,
as `{\it total} derivative with respect to
a given variable'.  The terminology suggested by Brownstein~\cite{brown} is
`the {\it whole}-partial derivative'.

Let us study the case when we deal with explicit and implicit dependencies 
$f ({\bf p}, E ({\bf p}))$. It is well known that the energy in the
relativism is connected with the 3-momentum as $E=\pm \sqrt{{\bf p}^2 +m^2}$%
; the unit system $c=\hbar=1$ is used. In other words, we must choose the
3-dimensional hyperboloid from the entire Minkowski space and the energy is 
{\it not} an independent quantity anymore. Let us calculate the commutator
of the whole derivative $\hat\partial /\hat\partial E$ and $\hat\partial / 
\hat\partial p_i$.
In order to make distinction between differentiating the explicit function
and that which contains both explicit and implicit dependencies, the `whole
partial derivative' may be denoted as $\hat\partial$.
In the general case one has 
\begin{equation}
{\frac{\hat\partial f ({\bf p}, E({\bf p})) }{\hat\partial p_i}} \equiv {%
\frac{\partial f ({\bf p}, E({\bf p})) }{\partial p_i}} + {\frac{\partial f (%
{\bf p}, E({\bf p})) }{\partial E}} {\frac{\partial E}{\partial p_i}}\, .
\end{equation}
Applying this rule, we surprisingly find 
\begin{eqnarray}
&&[{\frac{\hat\partial }{\hat\partial p_i}},{\frac{\hat\partial }{\hat%
\partial E}}] f ({\bf p},E ({\bf p})) = {\frac{\hat\partial }{\hat\partial
p_i}} {\frac{\partial f }{\partial E}} -{\frac{\partial }{\partial E}} ({%
\frac{\partial f}{\partial p_i}} +{\frac{\partial f}{\partial E}}{\frac{%
\partial E}{\partial p_i}}) =  \nonumber \\
&=& {\frac{\partial^2 f }{\partial E\partial p_i}} + {\frac{\partial^2 f}{%
\partial E^2}}{\frac{\partial E}{\partial p_i}} - {\frac{\partial^2 f }{%
\partial p_i \partial E}} - {\frac{\partial^2 f}{\partial E^2}}{\frac{%
\partial E}{\partial p_i}}- {\frac{\partial f }{\partial E}} {\frac{\partial%
}{\partial E}}({\frac{\partial E}{\partial p_i}})\,.  \label{com}
\end{eqnarray}
So, if $E=\pm \sqrt{m^2+{\bf p}^2}$ 
and one uses the generally-accepted 
representation form of $\partial E/\partial p_i
=  p_i/E$,
one has that the expression (\ref{com})
appears to be equal to $(p_i/E^2) {\frac{\partial f({\bf p}, E ({\bf p}))}{%
\partial E}}$. Within the choice of the normalization the coefficient may be related to 
the longitudinal electric field in the helicity basis (the electric/magnetic
fields can be derived from the 4-potentials which have been presented in~ 
\cite{hb}). On the other hand, the commutator 
\begin{equation}
[{\frac{\hat\partial}{\hat\partial p_i}}, {\frac{\hat\partial}{\hat\partial
p_j}}] f ({\bf p},E ({\bf p})) = {\frac{1}{\vert E\vert^3}} {\frac{%
\partial f({\bf p}, E ({\bf p}))}{\partial E}} [p_i, p_j]\,.
\end{equation}
This should also not  be zero according to
Feynman and Dyson~\cite{FD}. They postulated that the velocity (or, of course, the 3-momentum)
commutator is equal to $[p_i,p_j]\sim i\hbar\epsilon_{ijk} B^k$, i.e., to
the magnetic field.
In fact, if we put in the corespondence to the momenta their
quantum-mechanical operators (of course, with the appropriate clarification $%
\partial \rightarrow \hat\partial$), we obtain again that, in general, the
derivatives do {\it not} commute $[{\frac{\hat\partial}{\hat\partial x_\mu}}%
, {\frac{\hat\partial}{\hat\partial x_\nu}}] \neq 0$.

Furthermore, since the energy derivative corresponds to the operator of time
and the $i$-component momentum derivative, to $\hat x_i$, we put forward the
following Ansatz in the momentum representation: 
\begin{equation}
[\hat x^\mu, \hat x^\nu] = \omega ({\bf p}, E({\bf p})) \,
F^{\mu\nu}_{\vert\vert}{\frac{\partial }{\partial E}}\,,
\end{equation}
with some weight function $\omega$ being different for different choices of
the antisymmetric tensor spin basis. In the modern literature the relation 
$[x^\mu, x^\nu]\sim F^{\mu\nu}$, the electromagnetic tensor, is frequently 
used~\cite{elmnoncom}.
However, the idea of the broken Lorentz invariance by this
method~\cite{prl} concurs with the idea of the {\it fundamental length}, first
introduced by V. G. Kadyshevsky~\cite{kadysh} on the basis of old papers by
M. Markov. Both ideas and corresponding theories are extensively discussed,
e.g.~\cite{amelino}. In my opinion, the main question is: what is the space
scale, when the relativity theory becomes incorrect.

\section*{Conclusions}

We found that the commutator of two derivatives may be {\it not} equal to
zero. As a consequence, for instance, the question arises, if the derivative 
$\hat\partial^2 f/\hat\partial p^\nu\hat\partial p^\mu$ is equal to the
derivative $\hat\partial^2 f/\hat\partial p^\mu\hat\partial p^\nu$ in all
cases?\footnote{%
The same question can be put forward when we have differentiation with
respect to the coordinates too, that may have impact on the correct
calculations of the problem of accelerated charge in classical
electrodynamics.} The presented consideration permits us to provide some
bases for non-commutative field theories and induces us to look for further
development of the classical analysis in order to provide a rigorous
mathematical basis for operations with functions which have both explicit
and implicit dependencies.

I am grateful to participants of conferences where this idea has been discussed~\cite{dvoeglazovaps}.
I greatly apreciate the referee clear reports.
I am grateful to R. Keys for his help with the French translation.

\end{thebibliography}
\end{document}